# Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting

### v2

@@ Appendix A Discussions
Appendix B More Related Work
B.1 Interactive Vision Models.
Interactive vision models are designed for tasks based on user inputs. As one of the fundamental tasks in computer vision, they are extensively researched with numerous breakthroughs [26, 6, 33, 60, 14]. Particularly, the advent of the Segment Anything Model (SAM) [24] has sparked a surge of progress, with applications spanning reconstruction [49], detection [47, 63], segmentation [71], image editing [13], and more. Compared to existing models, which typically rely on prompts such as clicks, boxes, or scribbles, in this work, we study visual prompts, a new prompt referring to the actual images of objects with arbitrary poses and styles. Visual prompts enable continuous detection and tracking of target objects, facilitating the immediate correction of failed detections at test time. This represents a departure from traditional prompt types, offering more natural and dynamic interactions with the visual content.
B.2 Test-time Adaptation
Test-time adaptation aims to improve model performance on test data by adapting the model to test samples, even in the presence of data shifts [69, 38, 39, 48, 53, 34]. However, this area remains relatively unexplored in the context of 3D detection. Recent work, such as MonoTTA [29], has explored domain adaptation for 3D detection by fine-tuning models on batch streaming data under different weather conditions. In contrast, online 3D detection does not modify the original model parameters. Instead, it leverages human feedback to achieve instance-level error correction, thereby enabling domain adaptation without compromising the model’s performance on the original domain.
Appendix C Implementation Details
C.1 Details of TTC System
Figure 8: Details of the TTC system.
This section elaborates on the detailed workflow of the TTC system. As shown in Figure 8, during online inference, once the TTC-3D Detector fails to recognize an object, users can click on the missing object within the image. Based on the user-provided click, the TTC-3D Detector identifies the corresponding 2D boxes, crops the relevant areas to obtain visual prompt patches, and subsequently updates the visual prompt buffer. In later frames of inference, the TTC-3D Detector applies these stored visual prompts to continuously detect and track previously missing 3D objects, achieving instant error correction during test-time and constantly enhancing the offline-trained 3D detectors after being deployed.
C.2 Visual Prompt Alignment
Figure 9: Concrete mechanism of visual prompt alignment. This figure illustrates monocular input. When multi-view images are employed, this alignment operation flattens the different views and still generates $N$ peak candidate positions.
We now delve into the implementation details of visual prompt alignment. As described in Figure 9, given the visual prompt $\mathcal{P}_{v}$, we first extract the visual prompt features $\mathcal{Z}_{v}$ using a lightweight encoder, e.g., ResNet18 [17] in our implementation. Then, we use two separate 2-layer perceptrons, each with 128 and 64 output channels, to align the channel dimensions of the prompt features and the input image features. This results in the aligned feature maps $\hat{\mathcal{Z}_{v}}\in\mathbb{R}^{M\times 64}$ and $\hat{\mathcal{Z}}\in\mathbb{R}^{KHW\times 64}$, where $M$ is the number of visual prompts, $K$ is the number of image views, $H$ and $W$ are the spatial dimensions of image features. Finally, we compute the cosine similarity ($\odot$ in Figure 9) between $\hat{\mathcal{Z}}_{v}$ and $\hat{\mathcal{Z}}$ to obtain the similarity map $\mathcal{S}\in\mathbb{R}^{M\times KHW}$, which encodes the alignments between the visual prompts and image features.
To solve the instance ambiguity issue, we propose to predict multiple spatial coordinates of different peak responses in the similarity map (highlighted by “orange box” in Figure 9). Specifically, we first multiply the image features $\mathcal{Z}$ with similarity map $\mathcal{S}$ ($\bigotimes$ in Figure 9), and then use a 2-layer perceptron with $N\times 2$ output channels to regress spatial coordinates of the $N$ peak responses. During training, we only backpropagate the localization loss for the positions that have the minimum loss with respect to the ground truth. In our implementation, $N$ equals 4. This multi-instance retrieval approach allows the TTC to handle cases where the visual prompt matches multiple candidate objects in the input image, improving the robustness of the online error correction.
C.3 Details of Entity Detection Score (EDS)
Table 4: Feasibility Analysis of the EDS Metric.
Method mAP (%) w.o. Rec (%) + Rec (%)
BEVFormer 36.5 49.1 35.8
TTC-BEVFormer 47.8 54.6 47.2
BEVFormer-V2 39.6 49.2 38.9
TTC-BEVFormer-V2 51.6 58.5 51.0
EDS is a class-agnostic version of nuScenes Detection Score (NDS), but further prioritizes the localization quality of target objects. Inspired by [64], we improve the original NDS computation by multiplying the recall rate and the mean True Positive metrics, $\mathbb{TP}$, as well, as illustrated below:
$\text{EDS}\!=\!\frac{1}{6}[3\text{mAP}\!+\!\text{Recall}\!\times\!\sum_{\text{mTP}\in\mathbb{TP}}(1-\text{min}(1,\text{mTP}))],$ (2)
The intuition behind this is simple. The larger the recall rate is, the more predictions are involved in the statistics of mTP. Compared to simply setting a recall threshold [1], the multiplication adjusts the weight of mTP to EDS according to its comprehensiveness and thus brings a more informative quantitative result. In Table 4, we analyze the effectiveness of this metric. The comparisons between multiplying recall rate or not on various TTC-3D Detectors show that EDS, incorporating recall into its calculation, does not alter the original overall trend. Furthermore, it effectively highlights the superiority of detecting missed objects, demonstrating its validity.
Appendix D Ablation Studies
In this section, we conduct a series of ablation studies. We base our TTC system on MonoDETR for its simple and efficient monocular setting except for experiments in Section D.6, Section D.7 and Section D.9.
D.1 Effect of Components in Visual Prompt Alignment.
Table 5: Effect of visual prompt alignment.
Sim. Loss Loc. Loss mAP (%) EDS (%)
- - 32.6 31.9
$\surd$ - 39.4 38.5
$\surd$ $\surd$ 43.3 42.8
Visual prompt alignment aims to localize objects via visual prompts in input images. We now evaluate its components. Table 5 presents experiments to verify the core components of this alignment, including similarity loss (Focal and Dice loss) and localization loss (supervising visual prompt localization, $\mathcal{X}_{v}$). While the alignment can be implicitly learned by attention mechanisms in the transformer decoder, incorporating explicit similarity supervision brings improvements of 6.8% mAP and 6.6% EDS. Further utilizing the position loss and one-to-$N$ mapping (to address instance ambiguity) boosts the performance to 43.3% mAP and 42.8% EDS. These results prove this alignment operation is a critical component enabling the TTC detectors to effectively detect target objects via visual prompts.
D.2 Instance Ambiguity in Visual Prompt Alignment
Table 6: Effect of the number of predicted positions $N$ of visual prompt alignment.
No. of Position Predictions mAP (%) $\uparrow$ Recall (%) $\uparrow$
1 39.9 62.1
4 43.3 69.1
8 43.0 69.4
To solve the instance ambiguity issue, we propose to regress $N$ positions of each visual prompt when performing the visual prompt alignment. This retrieves all objects with similar visual contents. In this study, we validate the effectiveness of this design by conducting ablation studies on the number of $N$. As listed in Table 6, when the number of position prediction $N$ equals 1, which means a one-to-one mapping for each visual prompt, the mAP and recall rate are 39.9% and 62.1%, respectively. Then, if we increase the $N$ to 4, effectively a one-to-four mapping, we obtain an mAP of 43.3% and a recall rate of 69.1%. This represents a 7% improvement in the recall rate, demonstrating that instance ambiguity is an important challenge in visual prompt-based detection, and the proposed one-to-$N$ mapping solution effectively addresses this issue.
D.3 Effect of Visual Prompt Alignment on Instance Awareness
Figure 11: Visual prompt showcase for Table 7.
Group mAP (%) $\uparrow$ EDS (%) $\uparrow$
a. 20.9 27.5
b. 13.8 18.9
Despite demonstrating that visual prompt alignment can enhance system performance, we remain uncertain whether the alignment can distinguish different objects based on visual prompts for instance-level matching. To investigate, we conduct an additional experiment, based on a similar setting with Table 3 of the main paper [^3] For reference, we employ the TTC-MonoDETR trained on 20% vehicle annotations and freeze the prompt buffer with predefined web prompts when inference., but fix the visual prompt buffer with images of black sedans solely (Figure 11). As shown in Table 7, as “Car” objects in nuScenes contain various types and colors, solely using black sedans as visual prompts leads to an 8.6% EDS drop. This underscores that the alignment operation effectively differentiates objects based on visual prompts, achieving instance-level matching and detection.
D.4 Statistics of Visual Prompt Buffer during Online Inference.
We design the visual prompt buffer to store missed objects during inference with video stream and introduce a “dequeue” mechanism to prevent the buffer from growing indefinitely. In this ablation study, we analyze the dynamic buffer size, as well as the number of enqueued and dequeued in each frame, to illustrate the behavior of visual prompt buffer during the online operation of TTC.
As shown in Figure 10, the visual prompt buffer exhibits three distinct behaviors during online inference in each nuScenes video clip: increasing, steady, and decreasing. In initial frames, many traffic cones are queued into the buffer due to the poor performance of deployed offline 3D detectors on cone objects. The buffer size thus grows quickly in initial frames to store visual prompts of missed objects for online rectification (Frames #0 to #4). The buffer size then stabilizes as the online detector consistently detects and tracks all objects of interest (Frames #4 to #20). Further, as the ego vehicle drives out of the scene, many previously enqueued objects no longer exist and are thus removed from the buffer automatically (Frames #20 to #40). The buffer finally becomes empty as the scene changes.
This demonstrates the effectiveness of visual prompt buffer, which consistently stores missed objects during online inference and corrects online errors. This dynamic behavior, exhibiting increasing, steady, and decreasing phases, highlights itself to manage stored visual prompts for robust 3D object detection performance throughout the online inference with the balance between latency and accuracy.
D.5 Visual Prompts with User Pertubations
Figure 12: Robustness against feedback noise.
In Section 4.3, we have demonstrated the robustness of our system on visual prompts from diverse sources, styles, poses, scenes, and timestamps. Considering visual prompts during online inference are derived from user clicking, this will have positional deviations to the perfect 2D center of the target object. Thus, we design experiments with positional perturbations during test-time inference to analyze the impact. Expressly, we set the maximum translation ratios at 0%, 10%, 20%, 30%, and 40% to the ground-truth 2D centers, and input them to TTC system to evaluate the robustness against such disturbances. Figure 12 shows that the TTC-MonoDETR exhibits minor performance drops under increasing positional perturbations, demonstrating strong robustness to these disturbances and highlighting significant potential for real-world applications.
D.6 Robustness of TTC to Reduced Feedback Frequency
Table 8: mAP comparisons of different prompt collection frequencies. “N=0” means having 2D feedbacks of missed detections at every frame; “N=2” means less 2D feedbacks collected every 2 frames. TTC system demonstrates robust error correction performance even with reduced feedback frequency.
Exp. on Human Feedback collected with different frame interval $N$. N
0 2 4 6
TTC-BEVFormer 47.8 47.5 47.3 46.9
TTC-Sparse4Dv2 53.5 53.1 52.4 51.8
TTC-MV2D 51.0 50.7 50.2 49.6
Considering the practical challenges of deploying this system, where human feedback inevitably involves some delay, it is difficult to provide feedback on missed detections for every frame without additional aids. To address this, we design experiments to simulate such scenarios and investigate the robustness of visual prompts under reduced human feedback frequency. As shown in Table 8, we update the prompt buffer at varying frame intervals to mimic feedback provided at different time intervals. TTC demonstrates robust error correction performance even with reduced feedback frequency (from every frame to every 6 frames), showcasing its feasibility for real-world applications where real-time feedback may not always be practical.
D.7 Comparisons to Single Object Tracking (SOT) with 3D Detection Head
Method mAP (%) $\uparrow$ EDS (%) $\uparrow$
SOT + MV2D 43.0 43.2
TTC-MV2D 51.0 51.0
In this section, we compare our TTC system with a simple implementation of online error correction that integrates SOT and 3D detection. Due to the lack of related work on single object 3D tracking, we use the ATOM algorithm for 2D tracking [12] and the MV2D detector as the 3D object detection head [57]. MV2D can leverage the 2D tracking results obtained by ATOM to initialize the query, thereby achieving the error correction ability similar to TTC system. For implementation, we record the missed objects of MV2D and store their reference images along with 2D bounding boxes in the prompt buffer, which are inputs ATOM requires. We do not perform dequeue operations to keep tracking all missing objects by ATOM.
As shown in Table 9, the TTC system, empowered by the OA module, significantly outperforms the combination of SOT and the 3D detector head in error correction performance. This result highlights that simply combining SOT with a 3D detection head does not effectively achieve online error correction, further demonstrating the effectiveness of our TTC system.
D.8 Analysis of Model Size and Latency
Method LL3DA MonoDETR TTC-MonoDETR
#Params 118M 68M 80M
FPS (Hz) 0.42 11.1 9.1
We further analyze the model size and latency of TTC system, primarily comparing them with recent LLM-based promptable 3D detection methods [4, 22]. We argue that, for online promptable systems that are developed for real-world applications, latency can come from into two parts. The first one is the unavoidable delay caused by humans from observing the error to reacting to provide prompts. This is inherent to any online prompt-based approach. The other one is the delay associated with the inference speed of the system itself. As the first one cannot be controlled by the system design, we focus on the latter here.
As shown in Table 10, LLM-based promtable methods, like LL3DA [4] exhibit high latency that is inadequate for autonomous driving deployments. In contrast, our TTC system introduces only a little extra latency compared to its base detector, which meets the real-time inference requirements and thus can be applicable for online autonomous driving systems.
D.9 Comparison with Offline Fine-tuning using User Feedback
Table 11: Comparisons between the TTC-MV2D and MV2D fine-tuned with feedback 2D annotations. “N=0” means having 2D feedbacks at every frame; “N=2” means less 2D feedbacks collected every 2 frames. TTC system, though without any extra training, still outperforms the offline 2D fine-tuned MV2D, especially when annotations are limited (N=6).
Exp. on Human Feedback collected with different frame interval $N$. N
0 2 4 6
MV2D + Offline fine-tune 43.4 42.7 41.0 39.9
TTC-MV2D 51.0 50.7 50.2 49.6
In this section, we compare our TTC method with the approach that collects missing objects during inference and subsequently fine-tunes with the collected test-time 2D ground truth. This is another approach for utilizing test-time human feedback, though with delays in further model fine-tuning.
We select MV2D [57] as the baseline since it relies on 2D detection results for 3D object detection, thus allowing it to utilize 2D human feedback annotations to fine-tune. Additionally, we collect 2D feedback from various frame intervals, as users cannot provide feedback at every frame. For MV2D, we use all the 2D box annotations collected at specified intervals for fine-tuning. For the TTC-MV2D, we update the prompt buffer at these specified intervals.
As shown in Table 11, TTC-MV2D outperforms the MV2D model fine-tuned with 2D feedback. Notably, the performance of MV2D fine-tuned with human feedback from larger frame intervals, which means less frequent human feedback, declines significantly compared to models fine-tuned with feedback from every frame. In contrast, the TTC system allows the MV2D to perform effectively even with less frequent prompts of missed detections. This finding highlights the practicality of the TTC design in utilizing a single corresponding visual prompt for streaming data, as it is impractical for users to provide 2D prompts at each frame.
Appendix E Qualitative Results
@@
  • In Section E.1 and Section E.2, we fix the prompt buffer with visual prompts of either labeled or novel, unlabeled 3D objects from the nuScenes dataset to detect targets in the nuScenes images.
  • In Section E.3, we test the performance with prompt buffer containing visual prompts in styles differing from the training distribution, such as Lego.
  • In Section E.4, we visualize the similarity maps on out-of-domain images, including YouTube driving videos and Internet-sourced visual prompts, demonstrating the generalization of the visual prompt alignment and the effectiveness of our TTC in reducing driving risks in non-standard scenarios.
@@
E.1 In-domain Visual Prompts on nuScenes “Seen” Objects
This visualization focuses on the in-domain detection performance of the TTC system on the nuScenes dataset. We utilize visual prompts from labeled objects in the nuScenes dataset, and demonstrate the system’s ability to effectively detect and track these target objects across different frames, as shown in Figure 13 and Figure 14. The results illustrate that our TTC can accurately localize and consistently track the target objects of interest within the nuScenes scenarios, showcasing its effectiveness in handling in-domain visual prompts.
E.2 In-domain Visual Prompts on nuScenes “Unseen” Objects
This visualization focuses on the TTC’s ability to handle visual prompts of objects not labeled in the nuScenes dataset. As shown in Figure 15, Figure 16, and Figure 17, our method demonstrates its potential to detect and track novel, out-of-distribution objects with these unseen visual prompts.
The results illustrate the TTC system’s capability to go beyond the training distribution and effectively localize and track objects that were not part of the original labeled dataset. This showcases the versatility and generalization ability of the visual prompt-based framework, enabling the detection of previously unseen objects. These findings highlight the potential of the TTC system to continuously expand its object detection capabilities by incorporating user-provided visual prompts, even for objects that were not included in the initial training data.
E.3 Out-domain Visual Prompts on nuScenes “Seen” Objects
This visualization focuses on the TTC’s performance with visual prompts in styles different from the training distribution. As shown in Figure 18 and Figure 19, the model can effectively detect target objects using visual prompts in various views and styles that diverge from the original training data.
These results demonstrate the potential of the TTC system to be extended to customized online 3D detection scenarios, where users can provide arbitrary visual prompts to guide the detection of objects of interest. The model’s ability to handle prompts across diverse style domains highlights its flexibility and versatility, a key advantage for enabling user-centric, interactive 3D perception systems.
E.4 Out-domain Visual Prompts on Real-world Examples
@@
As shown in Figure 20 and Figure 21, our method can effectively localize non-expected animals with higher responses in the regions where animals located, even though it was trained solely on the nuScenes dataset. Figure 22, Figure 23, Figure 24, Figure 25 further present examples of detection in a real-world scenario containing both vehicles and animals. Our method can detect all objects with their visual prompts simultaneously.
This further exemplifies the strong generalization capability of our TTC system, underscoring its potential for effectively handling challenging, edge-case scenarios on roads. The ability to accurately detect and localize unexpected objects beyond the training distribution highlights the robustness of the proposed approach, a key requirement for reliable autonomous driving systems.
Appendix F License of Assets
The adopted nuScenes dataset [1] is distributed under a CC BY-NC-SA 4.0 license. We implement the model based on mmDet3D codebase [8], which is released under the Apache 2.0 license.
@@ Appendix F License of Assets
Figure 24: Visualizations on real world scenarios with out-of-domain visual prompts. This figure demonstrates a case of real-world 3D object detection. We provide a visual prompt from a separate image as the pre-defined prompt and visualize it at the beginning of the sequence. As described, our TTC system can successfully detect the ”Deer” object using this stylized deer prompt sourced from the internet. This example highlights the TTC’s capability to effectively leverage diverse, user-supplied visual prompts to accurately identify target objects, even when the prompts are not directly from the same scene.
Figure 25: Visualizations on real world scenarios with out-of-domain visual prompts. Another case demonstrates the use of an internet-sourced ”Deer” visual prompt to detect the deer in a different real-world scenario. As shown, our TTC system effectively detects the deer even when it is partially obscured by snow. This example further illustrates the robust performance of the TTC framework in accurately localizing target objects, even in challenging environmental conditions, by leveraging flexible visual prompts provided by users.

### v3

@@ Appendix A Discussions
Appendix B Additional Explanations
This section provides visual explanations and additional tables supporting analyses in the main paper.
Figure 11: Intrinsic mismatch across datasets. Equal-size image patches around cars at comparable 3D distance/shape from nuScenes (left) and Waymo (right). Despite similar geometry, the apparent 2D scale differs markedly due to different focal lengths and camera rigs, inducing depth bias and explaining why naive cross-dataset transfer of vision-only 3D detectors degrades severely. Adapted from [zheng2023cross].
Figure 12: Outcomes of visual prompts and their FP impact. Final predictions consist of dense-query results and visual-prompt results. Visual prompts produce few above-threshold candidates, and those candidates fall into five cases (i–v). After confidence filtering and NMS, only a small portion contributes to false positives (v), explaining why FP/frame remains near baseline.
B.1 Cross-dataset Transfer.
Figure˜11 visualizes the intrinsic mismatch across datasets, adapted from [zheng2023cross]. Equal-size image patches from nuScenes (left) and Waymo (right) show that objects at comparable 3D distance and shape appear at different 2D scales due to camera focal-length and rig disparities. This visualization supports the cross-dataset transfer analysis in Section˜4.4, explaining why naive cross-dataset transfer of vision-only 3D detectors leads to depth bias and severe degradation.
B.2 False-Positive Impact.
Figure˜12 illustrates five cases of visual prompt results. After confidence filtering and NMS, valid prompts mainly produce correct detections or re-detections, while invalid prompts rarely survive. The remaining above-threshold cases (i–v) explain why TTC reduces missed detections without materially increasing false positives, as discussed in FP Analysis in Section˜4.4. To corroborate the deployment-relevant regime, Table 15 reports FP/frame and FN/frame at a confidence threshold of $\tau{=}0.3$; changes in FP/frame are negligible and not statistically significant across backbones.
TABLE 15: False-alarm metrics at a deployment-oriented confidence threshold ($\tau{=}0.3$).
Method FP/frame@0.3 $\downarrow$ FN/frame@0.3 $\downarrow$
MonoDETR 1.5 2.5
TTC-MonoDETR 1.5 2.1
MV2D 1.3 2.9
TTC-MV2D 1.3 2.2
Sparse4Dv2 1.1 3.0
TTC-Sparse4Dv2 1.1 2.2
BEVFormer 1.2 2.9
TTC-BEVFormer 1.2 2.1
BEVFormerV2-t8 1.2 2.9
TTC-BEVFormerV2-t8 1.2 2.1
RayDN 1.3 2.9
TTC-RayDN 1.3 2.1
StreamPETR 1.3 2.8
TTC-StreamPETR 1.3 2.3
Appendix C Qualitative Results
@@
  • In Section˜C.1 and Section˜C.2, we fix the prompt buffer with visual prompts of either labeled or novel, unlabeled 3D objects from the nuScenes dataset to detect targets in the nuScenes images.
  • In Section˜C.3, we test the performance with prompt buffer containing visual prompts in styles differing from the training distribution, such as Lego.
  • In Section˜C.4, we visualize the similarity maps on out-of-domain images, including YouTube driving videos and Internet-sourced visual prompts, demonstrating the generalization of the visual prompt alignment and the effectiveness of our TTC in reducing driving risks in non-standard scenarios.
@@
C.1 In-domain Visual Prompts on nuScenes “Seen” Objects.
This visualization focuses on the in-domain detection performance of the TTC system on the nuScenes dataset. We utilize visual prompts from labeled objects in the nuScenes dataset, and demonstrate the system’s ability to effectively detect and track these target objects across different frames, as shown in Figure˜13 and Figure˜14.
C.2 In-domain Visual Prompts on nuScenes “Unseen” Objects
This visualization focuses on the TTC’s ability to handle visual prompts of objects not labeled in the nuScenes dataset. As shown in Figure˜15, Figure˜16, and Figure˜17, our method demonstrates its potential to detect and track novel, out-of-distribution objects with these unseen visual prompts.
C.3 Out-domain Visual Prompts on nuScenes “Seen” Objects
This visualization focuses on the TTC’s performance with visual prompts in styles different from the training distribution. As shown in Figure˜18 and Figure˜19, the model can effectively detect target objects using visual prompts in various views and styles that diverge from the original training data.
C.4 Out-domain Visual Prompts on Real-world Examples
@@
As shown in Figure˜20 and Figure˜21, our method can effectively localize non-expected animals with higher responses in the regions where animals located, even though it was trained solely on the nuScenes dataset. Figure˜22, Figure˜23, Figure˜24, Figure˜25 further present examples of detection in a real-world scenario containing both vehicles and animals. Our method can detect all objects with their visual prompts simultaneously.
C.5 License of Assets
The adopted nuScenes dataset [nuscenes2019] is distributed under a CC BY-NC-SA 4.0 license. We implement the model based on mmDet3D codebase [mmdet3d2020], which is released under the Apache 2.0 license.
@@ C.5 License of Assets
Figure 24: Visualizations on real world scenarios with out-of-domain visual prompts. This figure demonstrates a case of real-world 3D object detection. We provide a visual prompt from a separate image as the pre-defined prompt and visualize it at the beginning of the sequence. As described, our TTC system can successfully detect the "Deer" object using this stylized deer prompt sourced from the internet. This example highlights the TTC’s capability to effectively leverage diverse, user-supplied visual prompts to accurately identify target objects, even when the prompts are not directly from the same scene.
Figure 25: Visualizations on real world scenarios with out-of-domain visual prompts. Another case demonstrates the use of an internet-sourced "Deer" visual prompt to detect the deer in a different real-world scenario. As shown, our TTC system effectively detects the deer even when it is partially obscured by snow. This example further illustrates the robust performance of the TTC framework in accurately localizing target objects, even in challenging environmental conditions, by leveraging flexible visual prompts provided by users.